Fast Ltd. is a public company that prepares its consolidated financial statements in accordance with IFRS. Its net income in Year 2 was $215,000, and shareholders’ equity at December 31, Year 2, was $1,950,000.

Mr. Lombardi, the major shareholder, has made an offer to buy out the other shareholders, delist the company, and take it private. Thereafter, the company will report under ASPE. You have identified the following two areas in which Fast’s accounting principles differ between IFRS and ASPE.

1. Fast incurred research and development costs of $515,000 in Year 1. Thirty percent of these costs were related to development activities that met the criteria for capitalization as an intangible asset. The newly developed product was brought to market in January, Year 2 and is expected to generate sales revenue for 10 years.

2. Fast acquired equipment at the beginning of Year 1 at a cost of $130,000. The equipment has a five-year life with no expected residual value and is depreciated on a straight-line basis. At December 31, Year 1, Fast compiled the following information related to this equipment:

Expected future cash flows from use of the equipment $ 107,000
Present value of expected future cash flows from use of the equipment 90,000
Net realizable value 87,000

Required:

(a) Determine the amount at which Fast should report each of the following on its balance sheet at December 31, Year 2, using (1) IFRS and (2) ASPE. Ignore the possibility of any additional impairment or reversal of impairment loss at the end of Year 2. Assume that Fast wants to minimize net income. (Leave no cells blank - be certain to enter "0" wherever required. Omit $ sign in your response.)

(b) Prepare a reconciliation of net income for Year 2 and shareholders’ equity at December 31, Year 2, under IFRS to an ASPE basis. (Omit $ sign in your response.)

  

Net Income Year 2 under ASPE $
S/E @ Dec 31, Year 2 under ASPE $

*answers 340,050 and 2,075,050 are incorrect
Answer #1
Treatment of intangible assets
APSE IFRS
The criteria for development activities of internally
generated intangible assets is consistent between ASPE
and IFRS, except that ASPE allows an accounting policy
choice to either:
i) Expense such expenditures as incurred; or
ii) Capitalize such expenditures as an intangible
asset, provided the recognition criteria are met
All development phase expenditures must be capitalized
if they meet the recognition criteria.
As Fast Ltd. Wants to minimize it's income, it can expense off development expense(R&D) which has been capitalised under IFRS-38.
Reseach and development IFRS APSE
R&D Dec 31, Yr 2(refer note 1) $123,600 0
Note.1 Calculation of balance of R&D at end of Yr. 2
R&D value capitalised = 515,000*30% = $154,500
Less: Amortization of R&D for 2 years (154500/10)*2 $30,900
$123,600
Impairment treatment
The basis for recognizing an impairment loss are slightly different under ASPE and IFRS. ASPE compares the carry value with fair value for an asset group or reporting unit.
IFRS compares the carrying amount with the recoverable amount.
The recoverable amount is defined as the higher of the fair value less cost to sell or the value in use.
Recoverable amount of equipment under IFRS
Higher of
i) Fair value less cost to sell(i.e.) Net realizable value $87,000
ii) Value in use(i.e. present value of cash inflow)    $ 90,000
Hence, recoverable amount = $90,000
Carrying amount of Assets 31 Dec, Yr-1 130,000 - 26,000 = $104,000 (Depreciation Yr1 = (130,000-0)/5 = 26,000
IFRS APSE
Impairment loss Carrying amount vs. Recoverable amount Carrying amount vs fair value
Carrying amount $104,000 $104,000
Recoverable amount $90,000 $107,000 (under APSE, carrying amount is said to be not recoverable and impairment is said to exist if the carrying amount exceeds both the un-discounted cash flows plus the disposable value. Here the sum of un-discounted cash flows plus the disposable value is the recoverable amount.)
Impairment loss $14,000 0 Under APSE Test of impairment is
(Carrying Amount > Recoverable Amount (un-discounted cash flows + disposable value)
Carrying value after impairment at 31dec Yr.1 $90,000 $104,000
Less: Depreciation for Yr. 2 $22,500 $26,000 (90,000/4 = 22,500 depreciation under IFRS)
Equipment @ Dec. 31 Yr. 2 $67,500 $78,000
(b) Reconciation of Net income for Year 2
Net income as per IFRS $215,000
Add: Impariment loss not recognised under APSE $14,000
Less: Depreciation difference under year 2 $3,500
Add: Amortization of intangible assets under IFRS $15,450
Net income as per APSE $240,950
Reconciation of shareholders' equity for Year 2
Stockholders' equity as per IFRS $1,950,000
Less: Research and development expense expensed in year 1 under APSE -$154,500
Add: Amortization for year 1 and 2 of R&D under IFRS $15,450
Difference in profit as per IFRS and as per APSE $25,950
$1,836,900